11/26/99 EVE shift summary

ledovskoy@fnal.gov
Sat, 27 Nov 1999 00:18:31 -0600

Shift Summary for EVE shift, 26-NOV-99

On Shift: Herman W., Sasha L.

Hours of beam available at all intensities: 8.50
Hours of beam available at nominal intensity: 8.50
Hours of E799 data (special runs not included): 8.00 (see note)

note: 5.0h of E799 data + 3.0h of E799 data without TRDs

Hours of special runs (laser scans, muon runs...):

--------------------------------------------------------------
Accounting of lost time: access to reset CAEN module, TRD HV debugging

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift:

========================================================================
Accesses during shift: one access to recycle CAEN modules. TRD's required
resetting.

========================================================================
Overview of shift: Alarm filled shift. Difficulties withe the CAEN HV,
Otherwise relatively smooth data
taking. Note that spill 276 to spill 430 contains data
without active TRD's, for run 15064.

========================================================================
Description of Shift:


run spill time comment
------------------------------------------------------------------------
15064 15:30 Shift begins

17:05 After listening alarm every spill for an hour we
raise SA4/SEM alarm limit from 13.0E-7 to 13.1E-7

276 17:42 ktevana errors and TRD alarm goes off at the same
time. Several kfrend to make triggers going
282 17:50 Taking triggers again. TRDs are off. Difficulties
to restart CAEN server software.

Valeri, Mike A. are in control room and Leo over
a modem investigating CAEN problem. Theo and Ashkan
contacted for SDAQ consultations.

19:13 onsplit requests UPM039 in ktev2/tps3d2

412 20:45 Request an access to work on CAEN module.
Mike A., Valeri and Sasha L. go on access.

430 21:09 TRDs are up and happy, triggers are going again.
CAEN module has to be cycled.

22:00 notice that KTEV DAQ Monitoring screen does not
update the spill number. The tape thermometer shows
the same temperature also. Do ">reaper ktev" thing.

496 22:40 ktevana errors (missing FFFF from stream 5 cr 53)
497 back to normal

========================================================================
Follow up Notes on Previous Shift:

>Channels 1476,2893 had gain drift of ~ -5% during this shift. Continue
>monitoring...

-----------------------------------------------------
Time 1476 Gain Drift 2893 Gain Drift
wrt dbase (%) wrt dbase (%)
-----------------------------------------------------
16:05 -5.27528 -5.18342
16:40 -5.4224 -5.05887
17:10 -5.41609 -4.98893
21:30 -5.33203 -5.06165
22:25 -5.48262 -5.11441
23:50 -5.45981 -4.99819
-----------------------------------------------------

========================================================================
Items to Follow up on Next Shift:

muon run, next run and next shift
polarity flip, and keep eye on HCC